# Hybrid collective excitations in topological superconductor/ferromagnetic insulator heterostructures

### v3

@@ III.1 Excitation-induced first-order corrections to the electronic Green’s function
$\delta\check{g}=\delta\check{g}_{\omega,k}e^{i\left(\omega t+\bm{k}\cdot\bm{r}\right)}+\delta\check{g}_{-\omega,-k}e^{-i\left(\omega^{*}t+\bm{k}\cdot\bm{r}\right)}.$ (16)
For brevity here and below we denote the complex value of frequency $\omega+i\kappa$, including the excitation decay rate $\kappa$, simply as $\omega$. The magnon correction $\delta\Delta$ to the superconducting gap takes an analogous form:
@@
Here $\delta\hat{\Delta}_{\pm}=antidiag(\delta\Delta_{\omega,k},\delta\Delta^{*}_{-\omega,-k})$ and $\delta\hat{\Delta}_{\mp}=antidiag(\delta\Delta_{-\omega,-k},\delta\Delta^{*}_{\omega,k})$. Substituting Eqs. (16)-(17) into Eqs. (13)-(14) and taking into account that $h_{y}=(\delta h/2)e^{i\omega t+\bm{k}\cdot\bm{r}}$, for retarded and advanced components we obtain
$-i\hbar Dk^{2}\hat{g}_{0}^{R,A}(\epsilon-\frac{\omega}{2})\delta\hat{g}_{\omega,k}^{R,A}+iDk_{x}\frac{\delta h}{2v_{f}}[\hat{g}_{0}^{R,A}(\epsilon-\frac{\omega}{2})\tau_{z}\hat{g}_{0}^{R,A}(\epsilon+\frac{\omega}{2})-\tau_{z}]=$
$\left[\left(\epsilon\pm i\Gamma\right)\tau_{z}+i\hat{\Delta},\delta\hat{g}_{\omega,k}^{R,A}\right]-\{\frac{\omega}{2}\tau_{z},\delta\hat{g}_{\omega,k}^{R,A}\}+i\left(\delta\hat{\Delta}_{\pm}\hat{g}_{0}^{R,A}(\epsilon+\frac{\omega}{2})-\hat{g}_{0}^{R,A}(\epsilon-\frac{\omega}{2})\delta\hat{\Delta}_{\pm}\right)-\hat{C}_{\phi}^{R,A},$
@@
$-i\hbar Dk^{2}\hat{g}_{0}^{R}(\epsilon-\frac{\omega}{2})\delta\hat{g}^{K}_{\omega,k}-i\hbar Dk^{2}\hat{g}_{0}^{K}(\epsilon-\frac{\omega}{2})\delta\hat{g}_{\omega,k}^{A}+iDk_{x}\frac{\delta h}{2v_{f}}\hat{g}_{m}^{-}=$
$\left[\epsilon\tau_{z}+i\hat{\Delta},\delta\hat{g}^{K}_{\omega,k}\right]-\{\left(\frac{\omega}{2}-i\Gamma\right)\tau_{z},\delta\hat{g}^{K}_{\omega,k}\}+i\left(\delta\hat{\Delta}_{\pm}\hat{g}_{0}^{K}(\epsilon+\frac{\omega}{2})-\hat{g}_{0}^{K}(\epsilon-\frac{\omega}{2})\delta\hat{\Delta}_{\pm}\right)-\hat{C}_{\phi}^{K}+\hat{C}_{\Gamma},$
@@
The second necessary component $\delta\hat{g}_{-\omega,-\bm{k}}$ is obtained from Eqs. (18)-(19) by the substitution $(\omega,\bm{k})\rightarrow(-\omega^{*},-\bm{k})$. It is seen that the superconducting response to the magnon is linear with respect to $\delta\bm{h}$ if the magnon has a non-zero component of the wave vector $k_{x}$ along the direction of the equilibrium exchange field.
@@
$[\delta\hat{g}_{\omega,k}^{R,A}]_{ij}=a_{ij}^{R,A}\delta\Delta_{\omega,k}+b_{ij}^{R,A}\delta\Delta_{-\omega,-k}^{*}$
$+c_{ij}^{R,A}(\hat{\bm{k}}\times\delta\bm{h}_{\omega,\bm{k}})\cdot\hat{z}+d_{ij}^{R,A}\delta\phi_{\omega,k},$ (20)
@@
$[\delta\hat{g}_{\omega,k}^{K}]_{ij}=a_{ij}^{K}\delta\Delta_{\omega,k}+b_{ij}^{K}\delta\Delta_{-\omega,-k}^{*}$
$+c_{ij}^{K}(\hat{\bm{k}}\times\delta\bm{h}_{\omega,\bm{k}})\cdot\hat{z}+d_{ij}^{K}\delta\phi_{\omega,k},$ (21)
@@
$\delta\Delta_{\omega,k}=\frac{\lambda}{4i}\int_{-\epsilon_{c}}^{\epsilon_{c}}d\epsilon[\delta f_{\omega,k}^{K}]_{12}.$ (22)
@@ III.2 Self-consistent calculation of the electric potential
$\left(\frac{\partial^{2}}{\partial z^{2}}-k^{2}\right)\phi_{\omega,k}(z)=4\pi e^{2}\nu f_{\omega,k}\delta(z),$ (27)
@@
$\phi(\bm{r},t)=\phi_{\omega,k}(z)e^{i(\omega t+\bm{k}\cdot\bm{r})}.$ (28)
@@
$\phi_{\omega,\bm{k}}(z)=\begin{cases}Ce^{-kz},z>0\\ Ce^{kz},z<0\\ \end{cases},$ (29)
where $C$ is to be found from the boundary condition at $z=0$: $\partial\phi_{\omega,k}/\partial z|_{z=+0}-\partial\phi_{\omega,k}/\partial z|_{z=-0}=4\pi e^{2}\nu f_{\omega,k}$. Then the 2D scalar potential $\phi_{\omega,k}(z=0)\equiv\phi_{\omega,k}$ takes the form:
$\phi_{\omega,k}=-\frac{\hbar^{2}\omega_{p}^{2}}{4\pi(k\xi)(\pi k_{B}T_{c})\Delta_{0}}f_{\omega,k},$ (30)
@@
Using Eq. (26) we can express scalar potential $\delta\phi_{\omega,k}$ in terms of the excitation correction to the Keldysh Green’s function. The scalar potential can be written as
$\delta\phi_{\omega,k}=-\frac{\tilde{\omega}_{p}^{2}}{(1+n_{\delta\phi})\tilde{\omega}_{p}^{2}+2k\xi}\left[n_{\delta\Delta}\delta\Delta_{\omega,k}\right$.
$\left.+n_{\delta\Delta^{*}}\delta\Delta_{-\omega,-k}^{*}+n_{\delta h}(\hat{\bm{k}}\times\delta\bm{h}_{\omega,\bm{k}})\cdot\hat{z}\right$], (31)
@@
$[\delta\hat{g}_{\omega,k}^{R,A,K}]_{ij}=$ $\mathscr{A}_{ij}^{R,A,K}\delta\Delta_{\omega,k}+\mathscr{B}_{ij}^{R,A,K}\delta\Delta_{-\omega,-k}^{*}$
$+\mathscr{C}_{ij}^{R,A,K}(\hat{\bm{k}}\times\delta\bm{h}_{\omega,\bm{k}})\cdot\hat{z}$, (36)
@@ III.3 Magnetic part of the hybrid collective excitations
The dynamics of the spin wave in the ferromagnetic insulator is described by the LLG equation:
In this equation $\bm{H_{eff}}=Km_{x}\hat{x}$, where $K$ is the uniaxial anisotropy constant, $D_{m}$ is the magnon stiffness and $\alpha$ is the Gilbert damping parameter. The last term represents the spin torque, which describes the back action of the TS on the FI via the interface exchange interaction between the FI magnetization and electron spin polarization $\bm{s}$ in TS. The factor $1/d_{f}$, where $d_{f}$ is the thickness of the FI layer, comes from averaging of the interface term over the FI layer width. The electron spin polarization $\bm{s}$ in the helical metal can be calculated via the electric current in the TS [86, 73]
$\bm{s}=-\frac{1}{2ev_{f}}(\hat{z}\times\bm{J})$ (41)
@@
$\bm{J}=\bm{J}_{\omega,k}e^{i\left(\omega t+\bm{k}\cdot\bm{r}\right)}+\bm{J}_{-\omega,-k}e^{-i\left(\omega^{*}t+\bm{k}\cdot\bm{r}\right)},$ (43)
@@
$\bm{J}_{\omega,k}=\frac{\sigma}{16e}\int d\epsilon\Bigl{\{}i\bm{k}\left(g_{0}^{K}(\epsilon-\frac{\omega}{2})(\delta g^{A}_{11}+\delta g^{A}_{22})\right$.
@@
$J_{\omega,\bm{k}}^{x}=(\hat{\bm{k}}\cdot\hat{x})\left[J_{\delta\Delta}\delta\Delta_{\omega,\bm{k}}+J_{\delta\Delta^{*}}\delta\Delta^{*}_{-\omega,-\bm{k}}\right$.
$\left.+J_{2,(\omega,\bm{k})}[\hat{\bm{k}}\times\delta\bm{h}]\cdot\hat{z}\right]+J_{0,\omega}(\delta\bm{h}\times\hat{z})\cdot\hat{x}$. (45)
Coefficients $J_{\delta\Delta}$, $J_{\delta\Delta^{*}}$, $J_{0,\omega}$ and $J_{2,(\omega,\bm{k})}$ are found from Eqs. (44) and (III.2).
@@ III.4 Calculation of the spectrum of the hybrid collective modes
The hybrid excitations, which consist of the superconducting order parameter excitations in the TS and magnons in FI, are investigated in the basis $\hat{\Psi}=(\delta\Delta^{a},\delta\Delta^{p},\delta h_{y},\delta h_{z})^{T}$. Here the first two components $\delta\Delta^{a}_{\omega,\bm{k}}=[\delta\Delta_{\omega,\bm{k}}+\delta\Delta_{-\omega,-\bm{k}}^{*}]/2$ and $\delta\Delta^{p}_{\omega,\bm{k}}=[\delta\Delta_{\omega,\bm{k}}-\delta\Delta_{-\omega,-\bm{k}}^{*}]/2i$ represent the amplitude and phase modes of the superconducting order parameter, respectively. The second two components describe the magnetic part of the excitation, that is the magnon, via the relation $\delta h_{y,z}=-J_{ex}M_{s}\delta m_{y,z}/(2\gamma d_{s})$. The spectrum of the hybridized excitations is found from the combination of the self-consistency equation (22) for the OP and the linearized with respect to $\delta\bm{m}$ and $\bm{s}$ LLG equation (40). The resulting linear matrix equation for finding the eigenmodes of the FI/TS system can be written in the form:
@@
$\hat{M}_{hh}=\left(\begin{array}[]{cc}i\omega&(\omega_{b}+i\alpha\omega)\\ -(\omega_{b}+i\alpha\omega)+c_{s}\left[J_{0,\omega}+{\rm sgn}(k_{x})J_{2,(\omega,\bm{k})}\right]&i\omega\end{array}\right),$ (60)
@@ IV Spectrum of the hybrid collective excitations
$\delta\Delta_{\omega,\bm{k}}^{p,a}=F_{\omega,k}^{p(a)}(\hat{\bm{k}}\times\delta\bm{h}_{\omega,\bm{k}})\cdot\hat{z}$, (61)
@@
$F_{\omega,\bm{k}}^{a}$ $=\frac{-C^{K}_{12}(\omega,\bm{k})\left(A_{12}^{K*}(-\omega^{*},-\bm{k})-1-B_{12}^{K*}(-\omega^{*},-\bm{k})\right)+C^{K*}_{12}(-\omega^{*},-\bm{k})\left(A_{12}^{K}(\omega,\bm{k})-1-B_{12}^{K}(\omega,\bm{k})\right)}{i\cdot det\left[\hat{M}_{\Delta\Delta}\right]},$ (62)
$F_{\omega,\bm{k}}^{p}$ $=\frac{C^{K}_{12}(\omega,\bm{k})\left(A_{12}^{K*}(-\omega^{*},-\bm{k})-1+B_{12}^{K*}(-\omega^{*},-\bm{k})\right)+C^{K*}_{12}(-\omega^{*},-\bm{k})\left(A_{12}^{K}(\omega,\bm{k})-1+B_{12}^{K}(\omega,\bm{k})\right)}{det\left[\hat{M}_{\Delta\Delta}\right]},$ (63)

### v4

@@ III.1 Excitation-induced first-order corrections to the electronic Green’s function
$\delta\check{g}=\delta\check{g}_{\omega_{r},k}e^{i\left(\omega t+\bm{k}\cdot\bm{r}\right)}+\delta\check{g}_{-\omega_{r},-k}e^{-i\left(\omega^{*}t+\bm{k}\cdot\bm{r}\right)}.$ (24)
where $\omega\equiv\omega_{r}+i\kappa$ is the complex value of frequency, including the excitation decay rate $\kappa$. The magnon correction $\delta\Delta$ to the superconducting gap takes an analogous form:
@@
Here $\delta\hat{\Delta}_{\pm}=antidiag(\delta\Delta_{\omega_{r},k},\delta\Delta^{*}_{-\omega_{r},-k})$ and $\delta\hat{\Delta}_{\mp}=antidiag(\delta\Delta_{-\omega_{r},-k},\delta\Delta^{*}_{\omega_{r},k})$. Substituting Eqs. (24)-(25) into Eqs. (21)-(22) and taking into account that $h_{y}=(\delta h/2)e^{i\omega t+\bm{k}\cdot\bm{r}}$, for retarded and advanced components we obtain
$-i\hbar Dk^{2}\hat{g}_{0}^{R,A}(\epsilon-\frac{\omega}{2})\delta\hat{g}_{\omega_{r},k}^{R,A}+iDk_{x}\frac{\delta h}{2v_{f}}[\hat{g}_{0}^{R,A}(\epsilon-\frac{\omega}{2})\tau_{z}\hat{g}_{0}^{R,A}(\epsilon+\frac{\omega}{2})-\tau_{z}]=$
$\left[\left(\epsilon\pm i\Gamma\right)\tau_{z}+i\hat{\Delta},\delta\hat{g}_{\omega_{r},k}^{R,A}\right]-\{\frac{\omega}{2}\tau_{z},\delta\hat{g}_{\omega_{r},k}^{R,A}\}+i\left(\delta\hat{\Delta}_{\pm}\hat{g}_{0}^{R,A}(\epsilon+\frac{\omega}{2})-\hat{g}_{0}^{R,A}(\epsilon-\frac{\omega}{2})\delta\hat{\Delta}_{\pm}\right)-\hat{C}_{\phi}^{R,A},$
@@
$-i\hbar Dk^{2}\hat{g}_{0}^{R}(\epsilon-\frac{\omega}{2})\delta\hat{g}^{K}_{\omega_{r},k}-i\hbar Dk^{2}\hat{g}_{0}^{K}(\epsilon-\frac{\omega}{2})\delta\hat{g}_{\omega_{r},k}^{A}+iDk_{x}\frac{\delta h}{2v_{f}}\hat{g}_{m}^{-}=$
$\left[\epsilon\tau_{z}+i\hat{\Delta},\delta\hat{g}^{K}_{\omega_{r},k}\right]-\{\left(\frac{\omega}{2}-i\Gamma\right)\tau_{z},\delta\hat{g}^{K}_{\omega_{r},k}\}+i\left(\delta\hat{\Delta}_{\pm}\hat{g}_{0}^{K}(\epsilon+\frac{\omega}{2})-\hat{g}_{0}^{K}(\epsilon-\frac{\omega}{2})\delta\hat{\Delta}_{\pm}\right)-\hat{C}_{\phi}^{K}+\hat{C}_{\Gamma},$
@@
The second necessary component $\delta\hat{g}_{-\omega_{r},-\bm{k}}$ is obtained from Eqs. (26)-(27) by the substitution $(\omega,\bm{k})\rightarrow(-\omega^{*},-\bm{k})$. It is seen that the superconducting response to the magnon is linear with respect to $\delta\bm{h}$ if the magnon has a non-zero component of the wave vector $k_{x}$ along the direction of the equilibrium exchange field.
@@
$[\delta\hat{g}_{\omega_{r},k}^{R,A}]_{ij}=a_{ij}^{R,A}\delta\Delta_{\omega_{r},k}+b_{ij}^{R,A}\delta\Delta_{-\omega_{r},-k}^{*}$
$+c_{ij}^{R,A}(\hat{\bm{k}}\times\delta\bm{h}_{\omega_{r},\bm{k}})\cdot\hat{z}+d_{ij}^{R,A}\delta\phi_{\omega_{r},k},$ (28)
@@
$[\delta\hat{g}_{\omega_{r},k}^{K}]_{ij}=a_{ij}^{K}\delta\Delta_{\omega_{r},k}+b_{ij}^{K}\delta\Delta_{-\omega_{r},-k}^{*}$
$+c_{ij}^{K}(\hat{\bm{k}}\times\delta\bm{h}_{\omega_{r},\bm{k}})\cdot\hat{z}+d_{ij}^{K}\delta\phi_{\omega_{r},k},$ (29)
@@
$\delta\Delta_{\omega_{r},k}=\frac{\lambda}{4i}\int_{-\epsilon_{c}}^{\epsilon_{c}}d\epsilon[\delta f_{\omega_{r},k}^{K}]_{12}.$ (30)
@@ III.2 Self-consistent calculation of the electric potential
$\left(\frac{\partial^{2}}{\partial z^{2}}-k^{2}\right)\phi_{\omega_{r},k}(z)=4\pi e^{2}\nu f_{\omega_{r},k}\delta(z),$ (35)
@@
$\phi(\bm{r},t)=\phi_{\omega_{r},k}(z)e^{i(\omega t+\bm{k}\cdot\bm{r})}.$ (36)
@@
$\phi_{\omega_{r},\bm{k}}(z)=\begin{cases}Ce^{-kz},z>0\\ Ce^{kz},z<0\\ \end{cases},$ (37)
where $C$ is to be found from the boundary condition at $z=0$: $\partial\phi_{\omega_{r},k}/\partial z|_{z=+0}-\partial\phi_{\omega_{r},k}/\partial z|_{z=-0}=4\pi e^{2}\nu f_{\omega_{r},k}$. Then the 2D scalar potential $\phi_{\omega_{r},k}(z=0)\equiv\phi_{\omega_{r},k}$ takes the form:
$\phi_{\omega_{r},k}=-\frac{\hbar^{2}\omega_{p}^{2}}{4\pi(k\xi)(\pi k_{B}T_{c})\Delta_{0}}f_{\omega_{r},k},$ (38)
@@
Using Eq. (34) we can express scalar potential $\delta\phi_{\omega_{r},k}$ in terms of the excitation correction to the Keldysh Green’s function. The scalar potential can be written as
$\delta\phi_{\omega_{r},k}=-\frac{\tilde{\omega}_{p}^{2}}{(1+n_{\delta\phi})\tilde{\omega}_{p}^{2}+2k\xi}\left[n_{\delta\Delta}\delta\Delta_{\omega_{r},k}\right$.
$\left.+n_{\delta\Delta^{*}}\delta\Delta_{-\omega_{r},-k}^{*}+n_{\delta h}(\hat{\bm{k}}\times\delta\bm{h}_{\omega_{r},\bm{k}})\cdot\hat{z}\right$], (39)
@@
$[\delta\hat{g}_{\omega_{r},k}^{R,A,K}]_{ij}=$ $\mathscr{A}_{ij}^{R,A,K}\delta\Delta_{\omega_{r},k}+\mathscr{B}_{ij}^{R,A,K}\delta\Delta_{-\omega_{r},-k}^{*}$
$+\mathscr{C}_{ij}^{R,A,K}(\hat{\bm{k}}\times\delta\bm{h}_{\omega_{r},\bm{k}})\cdot\hat{z}$, (44)
@@ III.3 Magnetic part of the hybrid collective excitations
The dynamics of the spin wave in the ferromagnetic insulator is described by the LLG equation (8). The electron spin polarization $\bm{s}$ in the helical metal can be calculated via the electric current in the TS according to Eq. (7).
@@
$\bm{J}=\bm{J}_{\omega_{r},k}e^{i\left(\omega t+\bm{k}\cdot\bm{r}\right)}+\bm{J}_{-\omega_{r},-k}e^{-i\left(\omega^{*}t+\bm{k}\cdot\bm{r}\right)},$ (49)
@@
$\bm{J}_{\omega_{r},k}=\frac{\sigma}{16e}\int d\epsilon\Bigl\{i\bm{k}\left(g_{0}^{K}(\epsilon-\frac{\omega}{2})(\delta g^{A}_{11}+\delta g^{A}_{22})\right$.
@@
$J_{\omega_{r},\bm{k}}^{x}=(\hat{\bm{k}}\cdot\hat{x})\left[J_{\delta\Delta}\delta\Delta_{\omega_{r},\bm{k}}+J_{\delta\Delta^{*}}\delta\Delta^{*}_{-\omega_{r},-\bm{k}}\right$.
$\left.+J_{2,(\omega_{r},\bm{k})}[\hat{\bm{k}}\times\delta\bm{h}_{\omega_{r},\bm{k}}]\cdot\hat{z}\right]+J_{0,\omega_{r}}(\delta\bm{h}_{\omega_{r},\bm{k}}\times\hat{z})\cdot\hat{x}$. (51)
Coefficients $J_{\delta\Delta}$, $J_{\delta\Delta^{*}}$, $J_{0,\omega_{r}}$ and $J_{2,(\omega_{r},\bm{k})}$ are found from Eqs. (50) and (III.2).
@@ III.4 Calculation of the spectrum of the hybrid collective modes
The hybrid excitations, which consist of the superconducting order parameter excitations in the TS and magnons in FI, are investigated in the basis $\hat{\Psi}=(\delta\Delta^{a},\delta\Delta^{p},\delta h_{y},\delta h_{z})^{T}$. Here the first two components $\delta\Delta^{a}_{\omega_{r},\bm{k}}=[\delta\Delta_{\omega_{r},\bm{k}}+\delta\Delta_{-\omega_{r},-\bm{k}}^{*}]/2$ and $\delta\Delta^{p}_{\omega_{r},\bm{k}}=[\delta\Delta_{\omega_{r},\bm{k}}-\delta\Delta_{-\omega_{r},-\bm{k}}^{*}]/2i$ represent the amplitude and phase modes of the superconducting order parameter, respectively. The second two components describe the magnetic part of the excitation, that is the magnon, via the relation $\delta h_{y,z}=-J_{ex}M_{s}\delta m_{y,z}/(2\gamma d_{s})$. The spectrum of the hybridized excitations is found from the combination of the self-consistency equation (30) for the OP and the linearized with respect to $\delta\bm{m}$ and $\bm{s}$ LLG equation (8). The resulting linear matrix equation for finding the eigenmodes of the FI/TS system can be written in the form:
@@
$\hat{M}_{hh}=\left(\begin{array}[]{cc}i\omega&(\omega_{b}+i\alpha\omega)\\ -(\omega_{b}+i\alpha\omega)+c_{s}\left[J_{0,\omega_{r}}+{\rm sgn}(k_{x})J_{2,(\omega_{r},\bm{k})}\right]&i\omega\end{array}\right),$ (66)
@@ IV Spectrum of the hybrid collective excitations
$\delta\Delta_{\omega_{r},\bm{k}}^{p,a}=F_{\omega_{r},k}^{p(a)}(\hat{\bm{k}}\times\delta\bm{h}_{\omega_{r},\bm{k}})\cdot\hat{z}$, (67)
@@
$F_{\omega_{r},\bm{k}}^{a}$ $=\frac{-C^{K}_{12}(\omega,\bm{k})\left(A_{12}^{K*}(-\omega^{*},-\bm{k})-1-B_{12}^{K*}(-\omega^{*},-\bm{k})\right)+C^{K*}_{12}(-\omega^{*},-\bm{k})\left(A_{12}^{K}(\omega,\bm{k})-1-B_{12}^{K}(\omega,\bm{k})\right)}{i\cdot det\left[\hat{M}_{\Delta\Delta}\right]},$ (68)
$F_{\omega_{r},\bm{k}}^{p}$ $=\frac{C^{K}_{12}(\omega,\bm{k})\left(A_{12}^{K*}(-\omega^{*},-\bm{k})-1+B_{12}^{K*}(-\omega^{*},-\bm{k})\right)+C^{K*}_{12}(-\omega^{*},-\bm{k})\left(A_{12}^{K}(\omega,\bm{k})-1+B_{12}^{K}(\omega,\bm{k})\right)}{det\left[\hat{M}_{\Delta\Delta}\right]},$ (69)